OpenStudy (anonymous):

10^x+4=1000 10^x+4=1000 @Mathematics

OpenStudy (anonymous):

logathirmic

OpenStudy (anonymous):

is the x+4 all in the exponent of not?

OpenStudy (anonymous):

*or not

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright one sec

OpenStudy (anonymous):

There's two ways to do this but I'll just show you the way I think is easier. So the equation is: \[10^{x+4}=1000\] First I would manipulate the right side of the equation (1000), so that it becomes an exponent with a base of 10: \[1000=10^{3}\] Now you have: \[10^{x+4}=10^{3}\] The reason we do that is so that now we can take the base-10 log of both sides to greatly simplify the equation. By definition: \[\log_{10} (10^n)=n\] So when we take the log of both sides we get: \[\log_{10}(10^{x+4})=\log_{10}(10^{3})\] \[x+4=3\] \[x+4-4=3-4=-1\] So the final answer is: \[x=-1\] Checking our answer: \[10^{-1+4}=10^{3}=1000 \rightarrow \sqrt{}\]

OpenStudy (anonymous):

Sorry its so long but I wanted to be sure I fully explained it